# State-space kinetic Ising model reveals task-dependent entropy flow in sparsely active nonequilibrium neuronal dynamics

### v2

@@ 1.6 Evaluation of log marginal likelihood functions
The Taylor expansion of $q(\boldsymbol{\theta}^{i}_{t})$ up to the second order yields
$q(\boldsymbol{\theta}^{i}_{t})=q(\boldsymbol{\theta}^{*})+\left.\frac{\partial{q(\boldsymbol{\theta}^{i}_{t})}}{\partial{\boldsymbol{\theta}^{i}_{t}}}\right|_{\boldsymbol{\theta}^{i}=\boldsymbol{\theta}^{*}}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{*})+(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{*})^{\top}\frac{1}{2!}\left.\frac{\partial}{\partial\boldsymbol{\theta}^{i}_{t}}\left(\frac{\partial q(\boldsymbol{\theta}^{i})}{\partial(\boldsymbol{\theta}^{i}_{t})^{\top}}\right)\right|_{\boldsymbol{\theta}^{i}_{t}=\boldsymbol{\theta}^{*}}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{*})$ (153)
@@
The value of $\boldsymbol{\theta}^{i}_{t}$ that maximizes the function $q(\boldsymbol{\theta}^{i}_{t})$ is the MAP estimate $\boldsymbol{\theta}^{i}_{t|t}$ of the filter density. Further, the quadratic term evaluated at the MAP estimate is given by the inverse of the filter covariance $\mathbf{W}^{i}_{t|t}$. Hence, at $\boldsymbol{\theta}^{i}_{t}=\boldsymbol{\theta}^{i}_{t|t}$, the Taylor expansion becomes
$q(\boldsymbol{\theta}^{i}_{t})=q(\boldsymbol{\theta}^{*})+(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{*})^{\top}\frac{1}{2!}(\mathbf{W}^{i}_{t|t})^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{*})$ (154)
Hence the marginal likelihood is approximated as
@@
$p(\mathbf{x}_{0:T}|\mathbf{w})$ $=p(\mathbf{x}_{0}|\mathbf{w})\int\prod_{t=1}^{T}\prod_{i=1}^{N}\frac{\exp[q(\boldsymbol{\theta}^{i}_{t|t})]}{\sqrt{|2\pi\mathbf{W}^{i}_{t|t-1}|}}d\boldsymbol{\theta}^{i}_{t}\left\{\exp\left[-(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t})^{\top}\frac{1}{2}[\mathbf{W}^{i}_{t|t}]^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t})\right]\right\}$
@@
$=p(\mathbf{x}_{0}|\mathbf{w})\prod_{t=1}^{T}\prod_{i=1}^{N}\exp[q(\boldsymbol{\theta}^{i}_{t|t})]\frac{\sqrt{|2\pi\mathbf{W}^{i}_{t|t}|}}{\sqrt{|2\pi\mathbf{W}^{i}_{t|t-1}|}}$
$\quad\times\frac{1}{\sqrt{|2\pi\mathbf{W}^{i}_{t|t}|}}\int d\boldsymbol{\theta}^{i}_{t}\left\{\exp\left[-(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t})^{\top}\frac{1}{2}[\mathbf{W}^{i}_{t|t}]^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t})\right]\right\}$
$=p(\mathbf{x}_{0}|\mathbf{w})\prod_{t=2}^{T}\prod_{i=1}^{N}\sqrt{\frac{|2\pi\mathbf{W}^{i}_{t|t}|}{|2\pi\mathbf{W}^{i}_{t|t-1}|}}\exp[q(\boldsymbol{\theta}^{i}_{t|t})]$ (155)
@@
$\log[p(\mathbf{x}_{0:T}|\mathbf{w})]=\log[p(\mathbf{x}_{0}|\mathbf{w})]+\sum_{t=2}^{T}\sum_{i=1}^{N}[\frac{1}{2}\log|\mathbf{W}^{i}_{t|t}|-\frac{1}{2}\log|\mathbf{W}^{i}_{t|t-1}|+q(\boldsymbol{\theta}^{i}_{t|t})].$ (156)
2 S2 text. An alternative calculation of the backward conditional entropy
Here we give an alternative approach to obtain the backward conditional entropy to the one given in Methods. The result gives an identical approximate solution.
@@
$p(\mathbf{x}_{t-2})$ $=Q(\mathbf{x}_{t-2}),$ (157)
$p(\mathbf{x}_{t}|\mathbf{x}_{t-1})$ $=Q(\mathbf{x}_{t}).$ (158)
@@
$=-\sum_{\mathbf{x}_{t-2}}\sum_{\mathbf{x}_{t-1}}p(\mathbf{x}_{t-1}|\mathbf{x}_{t-2})Q(\mathbf{x}_{t-2})\sum_{\mathbf{x}_{t}}Q(\mathbf{x}_{t})\sum_{i}\left[x_{i,t-1}h_{i,t}(\mathbf{x}_{t})-\psi(h_{i,t}(\mathbf{x}_{t}))\right]$
$=-\sum_{i}\sum_{x_{i,t-1}}\sum_{\mathbf{x}_{t-2}}p(x_{i,t-1}|\mathbf{x}_{t-2})Q(\mathbf{x}_{t-2})\sum_{\mathbf{x}_{t}}Q(\mathbf{x}_{t})\left[x_{i,t-1}h_{i,t}(\mathbf{x}_{t})-\psi(h_{i,t}(\mathbf{x}_{t}))\right]$ (159)
and define $\tilde{\phi}_{i}(x_{i,t})$ as
$\tilde{\phi}(x_{i,t-1})=\sum_{\mathbf{x}_{t}}Q(\mathbf{x}_{t})\left[x_{i,t-1}h_{i,t}(\mathbf{x}_{t})-\psi(h_{i,t}(\mathbf{x}_{t}))]\right],$ (160)
@@
$\gamma(h_{i,t})=x_{i,t-1}h_{i,t}-\psi(h_{i,t}),$ (161)
$\tilde{\phi}_{i}(x_{i,t})$ can be approximated as
$\tilde{\phi}(x_{i,t-1})$ $=\sum_{\mathbf{x}_{t}}Q\left(\mathbf{x}_{t}\right)\left[x_{i,t-1}h_{i,t}(\mathbf{x}_{t})-\psi(h_{i,t}(\mathbf{x}_{t}))]\right]$
$\approx\int\mathcal{D}_{z}\,\gamma(g_{i,t}+z\sqrt{\Delta_{i,t}}),$ (162)
@@
${\sigma}^{\rm backward}_{t}$ $=-\sum_{\mathbf{x}_{t-2}}\sum_{\mathbf{x}_{t-1}}p(\mathbf{x}_{t-1}|\mathbf{x}_{t-2})Q(\mathbf{x}_{t-2})\tilde{\phi}(x_{i,t-1})$
$=-\sum_{i}\sum_{x_{i,t-1}}\left(\sum_{\mathbf{x}_{t-2}}p(x_{i,t-1}|\mathbf{x}_{t-2})Q(\mathbf{x}_{t-2})\right)\tilde{\phi}(x_{i,t-1})$ (163)
@@
$1-m_{i,t}$ $=\sum_{\mathbf{x}_{t-1}}p(\mathbf{x}_{t}=0|\mathbf{x}_{t-1})p(\mathbf{x}_{t-1})\simeq\sum_{\mathbf{x}_{t-1}}p(\mathbf{x}_{t}=0|\mathbf{x}_{t-1})Q(\mathbf{x}_{t-1}).$ (164)
@@
${\sigma}^{\rm backward}_{t}$ $\simeq-\sum_{i}\left\{m_{i,t-1}\tilde{\phi}(x_{i,t-1}=1)+(1-m_{i,t-1})\tilde{\phi}(x_{i,t-1}=0)\right\},$
$\simeq-\sum{i}\left\{m_{i,t-1}(\tilde{\phi}(x_{i,t-1}=1)-\tilde{\phi}(x_{i,t-1}=0))+\tilde{\phi}(x_{i,t-1}=0)\right\}.$ (165)
Thus it can be obtained by computing the two Gaussian integral terms.
@@
${\sigma}^{\rm backward}_{t}$ $\simeq-\sum_{i}\left\{m_{i,t-1}(\tilde{\phi}(x_{i,t-1}=1)-\tilde{\phi}(x_{i,t-1}=0))+\tilde{\phi}(x_{i,t-1}=0)\right\}.$ (166)
@@
$\tilde{\phi}(x_{i,t-1}=1)-\tilde{\phi}(x_{i,t-1}=0)$ $=\sum_{\mathbf{x}_{t}}Q\left(\mathbf{x}_{t}\right)h_{i,t}(\mathbf{x}_{t}),$
$\tilde{\phi}(x_{i,t-1}=0)$ $=-\sum_{\mathbf{x}_{t}}Q\left(\mathbf{x}_{t}\right)\psi(h_{i,t}(\mathbf{x}_{t})),$ (167)
@@
${\sigma}^{\rm backward}_{t}$ $=-\sum_{i}\sum_{\mathbf{x}_{t}}Q(\mathbf{x}_{t})\left[m_{i,t-1}h_{i,t}(\mathbf{x}_{t})-\psi(h_{i,t}(\mathbf{x}_{t}))\right],$ (168)
@@
3 S3 text. Mean-field entropy flow under specific conditions
@@
$\approx\sum_{i}\int\mathcal{D}_{z}\,\left[-\chi\left(g_{i,t,t-1}+z\sqrt{\Delta_{i,t,t-1}}\right)+\phi_{i}\left(g_{i,t,t}+z\sqrt{\Delta_{i,t,t}}\right)\right],$ (169)
@@
$g_{i,t,s}$ $=\theta_{i,t}+\sum_{j}\theta_{ij,t}m_{j,s},$ (170)
$\Delta_{i,t,s}$ $=\sum_{j}\theta_{ij,t}^{2}m_{j,s}(1-m_{j,s}).$ (171)
Here $m_{i,t}$ is the mean-field activation rate of the $i$-th neuron at time $t$.
Using $r(h)=1/(1+e^{-h})$ and $\psi(h)=-\log(1-r(h))$, $\chi(h)$ and $\phi_{i}(h)$ are given as
@@ 3.1 Steady-state solution
Under the steady-state assumption ($m_{i,t}=m_{i,t-1}\equiv m_{i}$), we have $g_{i,t,t-1}=g_{i,t,t}\equiv g_{i}$ and $\Delta_{i,t,t-1}=\Delta_{i,t,t}\equiv\Delta_{i}$, making the inputs to $\chi$ and $\phi_{i}$ common for each neuron. Then, using Eqs. 172 and 173 with the common $h=g_{i}+z\sqrt{\Delta_{i}}$, we have
@@
$=\sum_{i}\int\mathcal{D}_{z}\,\left(r\left(g_{i}+z\sqrt{\Delta_{i}}\right)-m_{i}\right)\cdot z\sqrt{\Delta_{i}}$ (174)
@@
${\sigma}^{\rm flow}_{t}\approx\sum_{i}\Delta_{i}\left(\int\mathcal{D}_{z}\,r^{\prime}(g_{i}+z\sqrt{\Delta_{i}})\right),$ (175)
where $r^{\prime}(h)=r(h)(1-r(h))$. Since $\Delta_{i}\geq 0$ and $r^{\prime}(h)\geq 0$, the entropy flow is non-negative, which satisfies the requested property of the entropy flow at the steady state. However, while insightful, this form also reveals a key limitation of the approximation: the zero entropy flow is realized only at $\theta_{ij}=0$ (except for $r=0,1$). Consequently, it does not correctly reduce to zero for symmetric couplings, failing to incorporate the distinction between symmetric and asymmetric interactions fully.

### v3

@@ 1.6 Approximate log marginal likelihood function
The Taylor expansion of $q(\boldsymbol{\theta}^{i}_{t})$ around $\boldsymbol{\theta}^{*}$ up to the second order yields
$q(\boldsymbol{\theta}^{i}_{t})=q(\boldsymbol{\theta}^{*})+\left.\frac{\partial{q(\boldsymbol{\theta}^{i}_{t})}}{\partial{\boldsymbol{\theta}^{i}_{t}}}\right|_{\boldsymbol{\theta}_{t}^{i}=\boldsymbol{\theta}^{*}}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{*})+\frac{1}{2}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{*})^{\top}\left.\frac{\partial^{2}q(\boldsymbol{\theta}^{i})}{\partial\boldsymbol{\theta}^{i}_{t}\partial(\boldsymbol{\theta}^{i}_{t})^{\top}}\right|_{\boldsymbol{\theta}^{i}_{t}=\boldsymbol{\theta}^{*}}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{*}).$ (S1.71)
@@
The value of $\boldsymbol{\theta}^{i}_{t}$ that maximizes the function $q(\boldsymbol{\theta}^{i}_{t})$ is the MAP estimate $\boldsymbol{\theta}^{i}_{t|t}$ of the filter density. Further, the quadratic term evaluated at the MAP estimate is given by the negative inverse of the filter covariance $\mathbf{W}^{i}_{t|t}$. Hence, at $\boldsymbol{\theta}^{*}=\boldsymbol{\theta}^{i}_{t|t}$, the Taylor expansion becomes
$q(\boldsymbol{\theta}^{i}_{t})\simeq q(\boldsymbol{\theta}^{i}_{t|t})-\frac{1}{2}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t})^{\top}(\mathbf{W}^{i}_{t|t})^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t}).$ (S1.72)
With this quadratic approximation, the marginal likelihood is obtained as
@@
$p(\mathbf{x}_{0:T}|\mathbf{w})$ $\simeq p(\mathbf{x}_{0})\prod_{t=1}^{T}\prod_{i=1}^{N}\int d\boldsymbol{\theta}^{i}_{t}\frac{1}{\sqrt{|2\pi\mathbf{W}^{i}_{t|t-1}|}}\exp\left[q(\boldsymbol{\theta}^{i}_{t|t})-(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t})^{\top}\frac{1}{2}[\mathbf{W}^{i}_{t|t}]^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t})\right]$
@@
$=p(\mathbf{x}_{0})\prod_{t=1}^{T}\prod_{i=1}^{N}\exp[q(\boldsymbol{\theta}^{i}_{t|t})]\frac{\sqrt{|2\pi\mathbf{W}^{i}_{t|t}|}}{\sqrt{|2\pi\mathbf{W}^{i}_{t|t-1}|}}\frac{1}{\sqrt{|2\pi\mathbf{W}^{i}_{t|t}|}}\int d\boldsymbol{\theta}^{i}_{t}\exp\left[-(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t})^{\top}\frac{1}{2}[\mathbf{W}^{i}_{t|t}]^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t})\right]$
$=p(\mathbf{x}_{0})\prod_{t=1}^{T}\prod_{i=1}^{N}\sqrt{\frac{|2\pi\mathbf{W}^{i}_{t|t}|}{|2\pi\mathbf{W}^{i}_{t|t-1}|}}\exp[q(\boldsymbol{\theta}^{i}_{t|t})].$ (S1.73)
@@
$\log p(\mathbf{x}_{0:T}|\mathbf{w})\simeq\log p(\mathbf{x}_{0})+\sum_{t=1}^{T}\sum_{i=1}^{N}\left[\frac{1}{2}\log|\mathbf{W}^{i}_{t|t}|-\frac{1}{2}\log|\mathbf{W}^{i}_{t|t-1}|+q(\boldsymbol{\theta}^{i}_{t|t})\right].$ (S1.74)
Supplementary Note 2 An alternative calculation of the backward conditional entropy
Here, we give an alternative approach to obtaining the backward conditional entropy to the one given in Methods. The result gives an identical approximate solution.
@@
$p(\mathbf{x}_{t-2})$ $=Q(\mathbf{x}_{t-2}),$ (S2.75)
$p(\mathbf{x}_{t}|\mathbf{x}_{t-1})$ $=Q(\mathbf{x}_{t}),$ (S2.76)
@@
$\simeq-\sum_{\mathbf{x}_{t-2}}\sum_{\mathbf{x}_{t-1}}p(\mathbf{x}_{t-1}|\mathbf{x}_{t-2})Q(\mathbf{x}_{t-2})\sum_{\mathbf{x}_{t}}Q(\mathbf{x}_{t})\sum_{i}\left[x_{i,t-1}h_{i,t}(\mathbf{x}_{t})-\psi(h_{i,t}(\mathbf{x}_{t}))\right]$
$=-\sum_{i}\sum_{x_{i,t-1}}\sum_{\mathbf{x}_{t-2}}p(x_{i,t-1}|\mathbf{x}_{t-2})Q(\mathbf{x}_{t-2})\sum_{\mathbf{x}_{t}}Q(\mathbf{x}_{t})\left[x_{i,t-1}h_{i,t}(\mathbf{x}_{t})-\psi(h_{i,t}(\mathbf{x}_{t}))\right].$ (S2.77)
Let us define $\tilde{\phi}_{i,t}(x_{i,t})$ as
$\tilde{\phi}_{i,t}(x_{i,t-1})=\sum_{\mathbf{x}_{t}}Q(\mathbf{x}_{t})\left[x_{i,t-1}h_{i,t}(\mathbf{x}_{t})-\psi(h_{i,t}(\mathbf{x}_{t}))\right].$ (S2.78)
@@
$\gamma(h_{i,t})=x_{i,t-1}h_{i,t}-\psi(h_{i,t}),$ (S2.79)
we approximate $\tilde{\phi}_{i,t}(x_{i,t})$ as
$\tilde{\phi}_{i,t}(x_{i,t-1})\approx\int\mathcal{D}_{z}\,\gamma(g_{i,t}+z\sqrt{\Delta_{i,t}}),$ (S2.80)
@@
${\sigma}^{\rm backward}_{t}$ $=-\sum_{\mathbf{x}_{t-2}}\sum_{\mathbf{x}_{t-1}}p(\mathbf{x}_{t-1}|\mathbf{x}_{t-2})Q(\mathbf{x}_{t-2})\tilde{\phi}_{i,t}(x_{i,t-1})$
$=-\sum_{i}\sum_{x_{i,t-1}}\left(\sum_{\mathbf{x}_{t-2}}p(x_{i,t-1}|\mathbf{x}_{t-2})Q(\mathbf{x}_{t-2})\right)\tilde{\phi}_{i,t}(x_{i,t-1}).$ (S2.81)
@@
$1-m_{i,t}$ $=\sum_{\mathbf{x}_{t-1}}p(\mathbf{x}_{t}=0|\mathbf{x}_{t-1})p(\mathbf{x}_{t-1})\simeq\sum_{\mathbf{x}_{t-1}}p(\mathbf{x}_{t}=0|\mathbf{x}_{t-1})Q(\mathbf{x}_{t-1}).$ (S2.82)
@@
${\sigma}^{\rm backward}_{t}$ $\simeq-\sum_{i}\left\{m_{i,t-1}\tilde{\phi}_{i,t}(x_{i,t-1}=1)+(1-m_{i,t-1})\tilde{\phi}_{i,t}(x_{i,t-1}=0)\right\}.$ (S2.83)
Thus, it can be obtained by computing the two Gaussian integral terms.
@@
${\sigma}^{\rm backward}_{t}$ $\simeq-\sum_{i}\left\{m_{i,t-1}(\tilde{\phi}_{i,t}(x_{i,t-1}=1)-\tilde{\phi}_{i,t}(x_{i,t-1}=0))+\tilde{\phi}_{i,t}(x_{i,t-1}=0)\right\}.$ (S2.84)
@@
$\tilde{\phi}_{i,t}(x_{i,t-1}=1)-\tilde{\phi}_{i,t}(x_{i,t-1}=0)$ $=\sum_{\mathbf{x}_{t}}Q\left(\mathbf{x}_{t}\right)h_{i,t}(\mathbf{x}_{t}),$
$\tilde{\phi}_{i,t}(x_{i,t-1}=0)$ $=-\sum_{\mathbf{x}_{t}}Q\left(\mathbf{x}_{t}\right)\psi(h_{i,t}(\mathbf{x}_{t})),$ (S2.85)
@@
${\sigma}^{\rm backward}_{t}$ $=-\sum_{i}\sum_{\mathbf{x}_{t}}Q(\mathbf{x}_{t})\left[m_{i,t-1}h_{i,t}(\mathbf{x}_{t})-\psi(h_{i,t}(\mathbf{x}_{t}))\right],$ (S2.86)
@@
Supplementary Note 3 Mean-field entropy flow under specific conditions
@@
$\approx\sum_{i}\int\mathcal{D}_{z}\,\left[-\chi\left(g_{i,t,t-1}+z\sqrt{\Delta_{i,t,t-1}}\right)+\phi_{i,t}\left(g_{i,t,t}+z\sqrt{\Delta_{i,t,t}}\right)\right],$ (S3.87)
@@
$g_{i,t,s}$ $=\theta_{i,t}+\sum_{j}\theta_{ij,t}m_{j,s},$ (S3.88)
$\Delta_{i,t,s}$ $=\sum_{j}\theta_{ij,t}^{2}m_{j,s}(1-m_{j,s}).$ (S3.89)
Here $m_{j,s}$ is the mean-field activation rate of the $j$-th neuron at time $s$.
Using $r(h)=1/(1+e^{-h})$ and $\psi(h)=-\log(1-r(h))$, $\chi(h)$ and $\phi_{i,t}(h)$ are given as
@@ 3.1 Steady-state solution
Under the steady-state assumption ($m_{i,t}=m_{i,t-1}\equiv m_{i}$), we have $g_{i,t,t-1}=g_{i,t,t}\equiv g_{i}$ and $\Delta_{i,t,t-1}=\Delta_{i,t,t}\equiv\Delta_{i}$, making the inputs to $\chi$ and $\phi_{i,t}$ common for each neuron. Then, using Eqs. S3.90 and S3.91 with the common $h=g_{i}+z\sqrt{\Delta_{i}}$, we have
@@
$=\sum_{i}\int\mathcal{D}_{z}\,\left(r\left(g_{i}+z\sqrt{\Delta_{i}}\right)-m_{i}\right)\cdot z\sqrt{\Delta_{i}}.$ (S3.92)
@@
${\sigma}^{\rm flow}_{t}\approx\sum_{i}\Delta_{i}\left(\int\mathcal{D}_{z}\,r^{\prime}(g_{i}+z\sqrt{\Delta_{i}})\right),$ (S3.93)
where $r^{\prime}(h)=r(h)(1-r(h))$. Since $\Delta_{i}\geq 0$ and $r^{\prime}(h)\geq 0$, the entropy flow is non-negative, which satisfies the requested property of the entropy flow at the steady state. However, while insightful, this form also reveals a key limitation of the approximation: the zero entropy flow is realized only at $\theta_{ij}=0$ (except for $r=0,1$). Consequently, it does not correctly reduce to zero for symmetric couplings, failing to fully incorporate the distinction between symmetric and asymmetric interactions.